# Structure Formation with Dark Magnetohydrodynamics

### v1

@@ V Effects on Matter Power Spectrum
$\ddot{\delta}_{\boldsymbol{k}}+2H\dot{\delta}_{\boldsymbol{k}}+\left[\frac{c_{s}^{2}(\mu)k^{2}}{a^{2}}-4\pi G\bar{\rho}_{m}\right]\delta_{\boldsymbol{k}}=0,\qquad c_{s}^{2}(\mu)=c_{s,\parallel}^{2}\,\mu^{2}+c_{s,\perp}^{2}(1-\mu^{2}),\qquad\mu\equiv\hat{\boldsymbol{k}}\cdot\hat{\boldsymbol{b}}$, (32)
@@
While assuming that $L_{B}\gg\lambda_{J}$, the effective group velocity, $c_{s}^{2}$ from Eq. (32), acquires a $k$- and angle-dependent pressure term, via Eq. (18) and Eq. (24):
@@
For the purpose of comparing altered matter power spectra from alternative DM models to that of CDM, it is convenient to define the transfer function, given by:
@@
where $P(k)$ is the power spectrum under an alternative DM model, and $P_{\Lambda\rm CDM}(k)$ is the isotropic CDM matter power spectrum. In our model, to compare with the isotropic matter power spectrum, we need to integrate over the directionality parameter $\mu$. To do so, we expand the power spectrum in terms of Legendre polynomials, $\mathcal{L}_{\ell}(\mu)$,
@@
Thus, the $\ell=0$ multipole will be the isotropic matter power spectrum from our model. In Fig. 3, we plot several examples of $P_{0}(k,0)$ for various values of the spectral index $n$. We show the current data from Planck 2018 Planck:2018nkj; Chabanier:2019eai, DES Y1 DES_Y1, SDSS DR7 SDSS_DR7, eBOSS DR14 Abolfathi_2018, MW Satellites MW_forecast, as well as projected future measurements from CMB-HD lensing MacInnis_2025, and the HERA and EDGES 21-cm projections HERA; EDGES; Munoz:2019hjh. In Fig. 4 we plot the corresponding transfer function of each of the curves from Fig. 3. We note that, unlike warm DM or SIDM models, we do not see pure exponential suppression or dark acoustic oscillations (DAOs), but rather a smooth reduction/alteration in power, followed by a continued power-law fall-off, with DAOs occurring at much larger wavenumbers. In principle, there is a scale such that all power is suppressed, though we do not probe these scales.
To place constraints on magnetic fields in the secluded $U(1)_{D}$ gauge sector, we scan over a 2D parameter space of the magnetic field today at reference scale $\lambda=0.1\,h^{-1}\,\rm{Mpc}$, and the DM charge-to-mass ratio. At each point in the scan, we numerically solve Eq. (32) to generate $P_{0}(k,0)$ and calculate the $\chi^{2}$ statistic against the current and future power spectrum data given in Fig. 3. The resulting parameters that can be excluded at the 68%, 95%, and 99.73% confidence levels are shown in the solid pink, purple, and red colored regions, respectively. Along with these contours, we also show projected constraints from future observations given by the dashed lines - the color scheme has the same meaning.
In each plot, we also fix the spectral index, $n$, to several representative values. For each index, the region is constrained from above by limits from CMB anisotropies PhysRevD.61.043001, noting that the strongest constraints are for the scale invariant power spectrum, n=3n=-3. Simulations of dark plasma instabilities in the Bullet Cluster additionally constrain the DM charge-to-mass ratio $q_{\chi}/m_{\chi}<2\times 10^{-14}\,\rm{GeV}^{-1}$ Giffin. Lastly, the Weak Gravity Conjecture Arkani_Hamed_2007 mandates that theories of quantum gravity with a $U(1)_{D}$ gauge field must satisfy $q_{\chi}/m_{\chi}>1/M_{P}$, where $M_{P}$ is the reduced Planck mass. This additionally constrains the dark charge-to-mass ratio from below. We find that with current data, the dark magnetic fields that have not already been excluded by measurements of the CMB anisotropies (with the small exception of the n=2n=-2 case, see Fig. 5) do not significantly alter the linear matter power spectrum, and are therefore consistent with current cosmology. With future measurements at higher wavenumbers, we expect to be able to probe deviations from the CDM power spectrum.
@@
Figure 5: Here we show the resultant parameter space for scans using our effective group velocity Eq. (34) to grow the linear density modes via Eq. (32) for current (shaded regions) and future (dashed lines) measurements of the linear matter power spectrum shown in Fig. 3 (see text for the data details). The region shown in pink is excluded at the $1\sigma$ level, the purple is excluded at the $2\sigma$ level, and the red region is excluded by a minimum of $3\sigma$. We also show constraint regions from the CMB PhysRevD.61.043001 (blue), Bullet Cluster Giffin (green), and the Weak Gravity Conjecture (yellow) Arkani_Hamed_2007.
@@ VI Discussion and Conclusions
In this work, we show that in the presence of large-scale dark magnetic fields, DM halos charged under a secluded $U(1)_{D}$ gauge field experience magnetic pressure that delays gravitational collapse via the Jeans instability. For correlation lengths much larger than $\lambda_{J}$, we solve Eq. (32) to map the effects onto the linear matter power spectrum. Fig. 5 shows that current measurements of the linear matter power spectrum provide constraints weaker than constraints set by tensor modes of the CMB. However, we expect future observations to probe into unconstrained regions of parameter space, further closing the gap between the constraint set by Ref. Giffin and the Weak Gravity Conjecture.
@@
Previous literature has examined the interactions of millicharged DM with SM particles and electromagnetic fields. In particular, the evolution of millicharged DM in the presence of plasma instabilities has been used as a powerful tool to constrain its properties, as demonstrated in studies of supernova remnants Li_2020 and the Bullet Cluster Cruz_2023. Considering a similar framework in which millicharged DM couples solely to SM electromagnetic fields, the model used in this work can be extended to incorporate baryonic effects by utilizing a two-fluid MHD model that includes gravity. Though SM primordial fields are generally better understood in the literature, the two-fluid model is sufficiently complex that the system likely needs to be solved numerically. We intend to consider this scenario in future work.
Another future direction is the consideration of the evolution of anisotropic collapse and halo triaxiality due to dark plasma effects. As a result of the directionally-dependent alteration to the Jeans Length, halo triaxiality and its radial evolution can serve as a direct macroscopic tracer of microscopic dark sector plasma effects. The anisotropic screening mechanism developed here predicts coherent, orientation-dependent modifications to collapse along and across the local dark-field direction $\hat{\boldsymbol{b}}$, leading to systematic biases in intrinsic shape distributions. Relative to the CDM baseline, dark-plasma screening enhances the contrast between inner and outer halo shapes and induces characteristic axis twists that preserve memory of the underlying field geometry. This “radial memory” manifests as steeper inner–outer ellipticity gradients and alignment signatures that cannot be replicated by baryonic condensation or isotropic self-interaction models. Triaxiality therefore emerges as a sensitive diagnostic of anisotropic support in the dark sector, complementing traditional small-scale power-spectrum constraints.
@@ VII Acknowledgments
The research of P.G., S.P., and M.G.R. is supported, in part, by the U.S. Department of Energy (DOE) grant number DE-SC0010107. AP acknowledges support from the Princeton Center for Theoretical Science and from the DOE under Award Number DE-SC0007968.
@@ Appendix A Derivation of Generalized Ohm’s Law
$\partial_{t}(\rho_{+}\vec{U}_{+})+\nabla\cdot(\rho_{+}\vec{U}_{+}\vec{U}_{+})=\rho_{+}\frac{q_{\chi}}{m_{\chi}}\left(\vec{E}+\vec{U}_{+}\times\vec{B}\right)-\nabla\cdot\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{+}-\rho_{+}\nu_{\pm}(\vec{U}_{+}-\vec{U_{-}})$ (38)
$\partial_{t}(\rho_{-}\vec{U}_{-})+\nabla\cdot(\rho_{-}\vec{U}_{-}\vec{U}_{-})=-\rho_{-}\frac{q_{\chi}}{m_{\chi}}\left(\vec{E}+\vec{U}_{-}\times\vec{B}\right)-\nabla\cdot\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{-}-\rho_{-}\nu_{\pm}(\vec{U}_{-}-\vec{U_{+}}),$ (39)
@@
We now consider the scenario under study. Given that we expect the positively changed species to be in thermal equilibrium with the negatively charged species, $\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{-}=\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{+}$ and the last term in Eq. (40) vanishes. Next, we assume that the unperturbed plasma initially has no net current or bulk velocity. Hence, after linearizing Eq. (40), $\nabla\cdot(\vec{U}\vec{J}+\vec{J}\vec{U})$ will only contain second order terms which we neglect in this work. Lastly, due to the strong constraints already placed on the dark charge-to-mass ratio by Giffin, we expect $\nu_{\pm}\ll T^{-1}$, where $T$ is the timescale of gravitational collapse of the plasma. Hence, the resistive term of Eq. (40) is negligible compared to the $\partial_{t}\vec{J}$ term. Thus, we chose to impliment the following Ohm’s law for our study
@@ Appendix B Derivation of Dispersion Relation
In this appendix, we explicitly show the full derivation of the dispersion relation in Eq. (III). These calculations follow very similarly to those performed in reference Gliddon_1966. However, we consider here effects from non-ideal MHD conditions. In the presence of an external magnetic field, a self-gravitating plasma obeys the following system equations
@@
Here we have utilized our choice of Ohm’s Law discussed in Appendix A as well as the Chew-Goldberger-Low equation of state for cylindrically symmetric magnetized plasmas CGL:1956. As in typical MHD studies, we assume that the plasma remains net neutral during its evolution and neglect effects due to separation of charges by setting $\rho_{c}=0$. Additionally, we assume that the system is varying slowly enough in time such that displacement current $\partial\vec{E}/\partial t$, can be neglected when compared to $\vec{J}$. This allows $\vec{J}$ to be eliminated from the system of equations
$\nabla\cdot\vec{E}=0$ $\nabla\times\vec{E}=-\frac{\partial\vec{B}}{\partial t}$ (49)
$\nabla\cdot\vec{B}=0$ (50)
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{U})=0$ $\rho\frac{d\vec{U}}{dt}=(\nabla\times\vec{B})\times\vec{B}-\nabla\stackrel{{\scriptstyle\leftrightarrow}}{{P}}-\rho\nabla V$ (51)
$\frac{d}{dt}\left(\frac{p_{\perp}B^{2}}{\rho^{3}}\right)=0$ $\frac{d}{dt}\left(\frac{p_{\parallel}}{\rho B}\right)=0$ (52)
$\vec{E}+\vec{U}\times\vec{B}=\frac{1}{2}\frac{m_{\chi}^{2}}{q_{\chi}^{2}\rho}\frac{\partial}{\partial t}\nabla\times\vec{B}$ $\nabla^{2}V=4\pi G\rho$ (53)
@@
$\nabla\times\left(\vec{U}\times\vec{B}-\frac{1}{2}\frac{m_{\chi}^{2}}{q_{\chi}^{2}\rho}\frac{\partial}{\partial t}\nabla\times\vec{B}\right)=\frac{\partial\vec{B}}{\partial t}$ $\nabla^{2}V=4\pi G\rho$ (54)
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{U})=0$ $\rho\frac{d\vec{U}}{dt}=(\nabla\times\vec{B})\times\vec{B}-\nabla\stackrel{{\scriptstyle\leftrightarrow}}{{P}}-\rho\nabla V$ (55)
$\frac{d}{dt}\left(\frac{p_{\perp}B^{2}}{\rho^{3}}\right)=0$ $\frac{d}{dt}\left(\frac{p_{\parallel}}{\rho B}\right)=0$ (56)
Note that in the $q_{\chi}/m_{\chi}\to\infty$ limit, we obtain the same system of equations studied under ideal MHD conditions in reference Gliddon_1966.
@@
$\rho\to\rho_{0}+\delta\rho,\quad\vec{U}\to\delta\vec{U},\quad\stackrel{{\scriptstyle\leftrightarrow}}{{P}}\to\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{0}+\stackrel{{\scriptstyle\leftrightarrow}}{{\delta P}},\quad\vec{B}\to\vec{B}_{0}+\delta\vec{B},\quad V\to\delta V_{0}.$ (60)
@@
$(\nabla\times\delta\vec{B})\times\vec{B}_{0}=-B_{0}^{2}\beta[(k_{x}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{x})\hat{x}+(k_{y}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{y})\hat{y}$] (79)
@@
$\rho_{0}\omega^{2}\xi_{x}=-\left[(1+\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}\xi_{z}\right]k_{x}p_{0}\\ -B_{0}^{2}\beta(k_{x}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{x})+4\pi G\rho_{0}^{2}\frac{k_{x}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (81)
$\rho_{0}\omega^{2}\xi_{y}=-\left[(1+\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}\xi_{z}\right]k_{y}p_{0}\\ -B_{0}^{2}\beta(k_{y}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{y})+4\pi G\rho_{0}^{2}\frac{k_{y}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (82)
$\rho_{0}\omega^{2}\xi_{z}=-[(3-2\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+3(k_{z}\xi_{z})]k_{z}p_{0}+4\pi G\rho_{0}^{2}\frac{k_{z}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (83)
@@
$\xi_{\parallel}=\xi_{z},\quad\xi_{\perp}=(\xi_{x},\xi_{y}),\quad k_{\parallel}=k_{z},\quad k_{\perp}=(k_{x},k_{y}),\quad k_{\perp}^{2}=k_{x}^{2}+k_{y}^{2}$ (84)
We now multiply Eq. (81) by $k_{x}$ and Eq. (82) by $k_{y}$ and add the resulting equations together to yield
$\left[\rho\omega^{2}-(1+\beta)k_{\perp}^{2}p-\beta B^{2}k^{2}+4\pi G\rho^{2}\frac{k_{\perp}^{2}}{k^{2}}\right]k_{\perp}\cdot\xi_{\perp}=\left(k_{\perp}^{2}p-4\pi G\rho^{2}\frac{k_{\perp}^{2}}{k^{2}}\right)k_{\parallel}\xi_{\parallel}$ (85)
Eq. (83) yields
$\left(\rho\omega^{2}-3k_{\parallel}^{2}p+4\pi G\rho^{2}\frac{k_{\parallel}^{2}}{k^{2}}\right)k_{\parallel}\xi_{\parallel}=\left((3-2\chi)k_{\parallel}^{2}p-4\pi G\rho^{2}\frac{k_{\parallel}^{2}}{k^{2}}\right)k_{\perp}\cdot\xi_{\perp}$ (86)

### v2

@@ V Effects on Matter Power Spectrum
$\ddot{\delta}_{\boldsymbol{k}}+2H\dot{\delta}_{\boldsymbol{k}}+\left[\frac{c_{s}^{2}(\mu)k^{2}}{a^{2}}-4\pi G\bar{\rho}_{m}\right]\delta_{\boldsymbol{k}}=0,$ (32)
with
$c_{s}^{2}(\mu)=c_{s,\parallel}^{2}\,\mu^{2}+c_{s,\perp}^{2}(1-\mu^{2}),\qquad\mu\equiv\hat{\boldsymbol{k}}\cdot\hat{\boldsymbol{b}}$, (33)
@@
While assuming that the coherence length of the dark magnetic field is larger than the MHD turbulence scale defined in Eq. (27) ($L_{B}\gg\ell$), the effective group velocity, $c_{s}^{2}$ from Eq. (33), acquires a $k$- and angle-dependent pressure term, via Eq. (18) and Eq. (24):
@@
For the purpose of comparing altered matter power spectra from alternative DM models to that of CDM, it is convenient to define the transfer function, given by
@@
where $P(k)$ is the power spectrum under an alternative DM model, and $P_{\Lambda\rm CDM}(k)$ is the isotropic CDM matter power spectrum. Here, we only study the effects of the linear evolution of density perturbations in Eq. (32). Though we expect our model to additionally impact the nonlinear matter power spectrum, such effects would require dedicated simulation which we consider to be beyond the scope of this paper. In our model, to compare with the isotropic matter power spectrum, we need to integrate over the directionality parameter $\mu$. To do so, we expand the power spectrum in terms of Legendre polynomials, $\mathcal{L}_{\ell}(\mu)$,
@@
Thus, the $\ell=0$ multipole will be the isotropic matter power spectrum from our model. In Fig. 3, we plot several examples of $P_{0}(k,0)$ for various values of the spectral index $n$. We show the current data from Planck 2018 [5, 21], DES Y1 [26], SDSS DR7 [79], eBOSS DR14 [1], MW Satellites [36], as well as projected future measurements from CMB-HD lensing [64], and the HERA and EDGES 21-cm projections [30, 62, 72]. In Fig. 4 we plot the corresponding transfer function of each of the curves from Fig. 3. We note that, unlike warm DM or SIDM models, we do not see pure exponential suppression or dark acoustic oscillations (DAOs), but rather a smooth reduction/alteration in power, followed by a continued power-law fall-off, with DAOs occurring at much larger wave numbers. In principle, there is a scale such that all power is suppressed, though we do not probe these scales. Additionally, it has recently been shown that the solutions to Eq. (33) closely match those from state-of-the-art numerical MHD simulations for scales larger than the MHD turbulent scale, $\lambda\gg\ell$ [77].
To place constraints on magnetic fields in the secluded $U(1)_{D}$ gauge sector, we scan over a 2D parameter space of the magnetic field today at reference scale $\lambda=0.1\,h^{-1}\,\rm{Mpc}$, and the DM charge-to-mass ratio. At each point in the scan, we numerically solve Eq. (33) to generate $P_{0}(k,0)$ and calculate the $\chi^{2}$ statistic against the current and future power spectrum data given in Fig. 3. The resulting parameters that can be excluded at the 68%, 95%, and 99.73% confidence levels are shown in the solid pink, purple, and red colored regions, respectively. Along with these contours, we also show projected constraints from future observations given by the dashed lines—the color scheme has the same meaning.
In each plot, we also fix the spectral index, $n$, to several representative values. For each index, the region is constrained from above by limits from CMB anisotropies [34], noting that the strongest constraints are for the scale invariant power spectrum, n=3n=-3. Simulations of dark plasma instabilities in the Bullet Cluster additionally constrain the DM charge-to-mass ratio $q_{\chi}/m_{\chi}<2\times 10^{-14}\,\rm{GeV}^{-1}$ [31]. Lastly, the weak gravity conjecture [8] mandates that theories of quantum gravity with a $U(1)_{D}$ gauge field must satisfy $q_{\chi}/m_{\chi}>1/M_{P}$, where $M_{P}$ is the reduced Planck mass. This additionally constrains the dark charge-to-mass ratio from below. We find that with current data, the dark magnetic fields that have not already been excluded by measurements of the CMB anisotropies (with the small exception of the n=2n=-2 case, see Fig. 5) do not significantly alter the linear matter power spectrum, and are therefore consistent with current cosmology. With future measurements at higher wave numbers, we expect to be able to probe deviations from the CDM power spectrum.
@@ VI Discussion and Conclusions
In this work, we show that in the presence of large-scale dark magnetic fields, DM halos charged under a secluded $U(1)_{D}$ gauge field experience magnetic pressure that delays gravitational collapse via the Jeans instability. For correlation lengths on scales much larger than the turbulent scale, $\ell$, we solve Eq. (33) to map the effects onto the linear matter power spectrum. Figure 5 shows that current measurements of the linear matter power spectrum provide constraints weaker than constraints set by tensor modes of the CMB. However, we expect future observations to probe into unconstrained regions of parameter space, further closing the gap between the constraint set by Ref. [31] and the weak gravity conjecture.
@@
Previous literature has examined the interactions of millicharged DM with SM particles and electromagnetic fields. In particular, the evolution of millicharged DM in the presence of plasma instabilities has been used as a powerful tool to constrain its properties, as demonstrated in studies of supernova remnants [63] and the Bullet Cluster [24]. Considering a similar framework in which millicharged DM couples solely to SM electromagnetic fields, the model used in this work can be extended to incorporate baryonic effects by utilizing a two-fluid MHD model that includes gravity. Though SM primordial fields are generally better understood in the literature, the two-fluid model is sufficiently complex that the system likely needs to be solved numerically. We intend to consider this scenario in future work.
Another future direction is the consideration of the evolution of anisotropic collapse and halo triaxiality due to dark plasma effects. As a result of the directionally dependent alteration to the Jeans length, halo triaxiality and its radial evolution can serve as a direct macroscopic tracer of microscopic dark sector plasma effects. The anisotropic screening mechanism developed here predicts coherent, orientation-dependent modifications to collapse along and across the local dark-field direction $\hat{\boldsymbol{b}}$, leading to systematic biases in intrinsic shape distributions. Relative to the CDM baseline, dark-plasma screening enhances the contrast between inner and outer halo shapes and induces characteristic axis twists that preserve memory of the underlying field geometry. This “radial memory” manifests as steeper inner-outer ellipticity gradients and alignment signatures that cannot be replicated by baryonic condensation or isotropic self-interaction models. Triaxiality therefore emerges as a sensitive diagnostic of anisotropic support in the dark sector, complementing traditional small-scale power-spectrum constraints.
@@
Figure 5: Here we show the resultant parameter space for scans using our effective group velocity Eq. (35) to grow the linear density modes via Eq. (33) for current (shaded regions) and future (dashed lines) measurements of the linear matter power spectrum shown in Fig. 3 (see text for the data details). The region shown in pink is excluded at the $1\sigma$ level, the purple is excluded at the $2\sigma$ level, and the red region is excluded by a minimum of $3\sigma$. We also show constraint regions from the CMB [34] (blue), Bullet Cluster [31] (green), and the weak gravity conjecture (yellow) [8].
@@ VII Acknowledgments
The research of P.G., S.P., and M.G.R. is supported, in part, by the U.S. Department of Energy (DOE) Grant No. DE-SC0010107. The research of P.G. is additionally supported in part by the Achievement Rewards for College Scientists Foundation (ARCS) 2025-2026. A.P. acknowledges support from the Princeton Center for Theoretical Science and from the DOE under Award Number DE-SC0007968.
@@ Appendix A DERIVATION OF GENERALIZED OHM’S LAW
$\partial_{t}(\rho_{+}\vec{U}_{+})+\nabla\cdot(\rho_{+}\vec{U}_{+}\vec{U}_{+})=\rho_{+}\frac{q_{\chi}}{m_{\chi}}\left(\vec{E}+\vec{U}_{+}\times\vec{B}\right)$
$\qquad\qquad\qquad\qquad-\nabla\cdot\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{+}-\rho_{+}\nu_{\pm}(\vec{U}_{+}-\vec{U_{-}}),$ (39)
$\partial_{t}(\rho_{-}\vec{U}_{-})+\nabla\cdot(\rho_{-}\vec{U}_{-}\vec{U}_{-})=-\rho_{-}\frac{q_{\chi}}{m_{\chi}}\left(\vec{E}+\vec{U}_{-}\times\vec{B}\right)$
$\qquad\qquad\qquad\qquad-\nabla\cdot\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{-}-\rho_{-}\nu_{\pm}(\vec{U}_{-}-\vec{U_{+}}),$ (40)
@@
We now consider the scenario under study. Given that we expect the positively changed species to be in thermal equilibrium with the negatively charged species, $\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{-}=\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{+}$ and the last term in Eq. (41) vanishes. Next, we assume that the unperturbed plasma initially has no net current or bulk velocity. Hence, after linearizing Eq. (41), $\nabla\cdot(\vec{U}\vec{J}+\vec{J}\vec{U})$ will only contain second order terms which we neglect in this work. Lastly, due to the strong constraints already placed on the dark charge-to-mass ratio by [31], we expect $\nu_{\pm}\ll T^{-1}$, where $T$ is the timescale of gravitational collapse of the plasma. Hence, the resistive term of Eq. (41) is negligible compared to the $\partial_{t}\vec{J}$ term. Thus, we chose to implement the following Ohm’s law for our study
@@ Appendix B DERIVATION OF DISPERSION RELATION
In this Appendix, we explicitly show the full derivation of the dispersion relation in Eq. (III). These calculations follow very similarly to those performed in Ref. [43]. However, we consider here effects from nonideal MHD conditions. In the presence of an external magnetic field, a self-gravitating plasma obeys the following system equations
@@
Here we have utilized our choice of Ohm’s Law discussed in Appendix A as well as the Chew-Goldberger-Low equation of state for cylindrically symmetric magnetized plasmas [22]. As in typical MHD studies, we assume that the plasma remains net neutral during its evolution and neglect effects due to separation of charges by setting $\rho_{c}=0$. Additionally, we assume that the system is varying slowly enough in time such that displacement current $\partial\vec{E}/\partial t$, can be neglected when compared to $\vec{J}$. This allows $\vec{J}$ to be eliminated from the system of equations
$\nabla\cdot\vec{E}=0,\qquad\nabla\times\vec{E}=-\frac{\partial\vec{B}}{\partial t},$ (50)
$\nabla\cdot\vec{B}=0,\qquad\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{U})=0,$ (51)
$\rho\frac{d\vec{U}}{dt}=(\nabla\times\vec{B})\times\vec{B}-\nabla\stackrel{{\scriptstyle\leftrightarrow}}{{P}}-\rho\nabla V,$ (52)
$\frac{d}{dt}\left(\frac{p_{\perp}B^{2}}{\rho^{3}}\right)=0,\qquad\frac{d}{dt}\left(\frac{p_{\parallel}}{\rho B}\right)=0,$ (53)
$\vec{E}+\vec{U}\times\vec{B}=\frac{1}{2}\frac{m_{\chi}^{2}}{q_{\chi}^{2}\rho}\frac{\partial}{\partial t}\nabla\times\vec{B},\qquad\nabla^{2}V=4\pi G\rho,$ (54)
@@
$\nabla\times\left(\vec{U}\times\vec{B}-\frac{1}{2}\frac{m_{\chi}^{2}}{q_{\chi}^{2}\rho}\frac{\partial}{\partial t}\nabla\times\vec{B}\right)=\frac{\partial\vec{B}}{\partial t}$ (55)
$\nabla^{2}V=4\pi G\rho,\qquad\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{U})=0,$ (56)
$\rho\frac{d\vec{U}}{dt}=(\nabla\times\vec{B})\times\vec{B}-\nabla\stackrel{{\scriptstyle\leftrightarrow}}{{P}}-\rho\nabla V,$ (57)
$\frac{d}{dt}\left(\frac{p_{\perp}B^{2}}{\rho^{3}}\right)=0,\qquad\frac{d}{dt}\left(\frac{p_{\parallel}}{\rho B}\right)=0.$ (58)
Note that in the $q_{\chi}/m_{\chi}\to\infty$ limit, we obtain the same system of equations studied under ideal MHD conditions in Ref. [43].
@@
$\rho\to\rho_{0}+\delta\rho,\quad\vec{U}\to\delta\vec{U},\quad\stackrel{{\scriptstyle\leftrightarrow}}{{P}}\to\stackrel{{\scriptstyle\leftrightarrow}}{{P}}_{0}+\stackrel{{\scriptstyle\leftrightarrow}}{{\delta P}},$ (61)
$\quad\vec{B}\to\vec{B}_{0}+\delta\vec{B},\quad V\to\delta V_{0}.$ (62)
@@
$(\nabla\times\delta\vec{B})\times\vec{B}_{0}=$ $-B_{0}^{2}\beta[(k_{x}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{x})\hat{x}$
$+(k_{y}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{y})\hat{y}$] (81)
@@
$-\rho_{0}\omega^{2}\xi_{x}=$ $-\left[(1+\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}\xi_{z}\right]k_{x}p_{0}$
$-B_{0}^{2}\beta(k_{x}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{x})$
$+4\pi G\rho_{0}^{2}\frac{k_{x}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (83)
$-\rho_{0}\omega^{2}\xi_{y}=$ $-\left[(1+\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}\xi_{z}\right]k_{y}p_{0}$
$-B_{0}^{2}\beta(k_{y}(k_{x}\xi_{x}+k_{y}\xi_{y})+k_{z}^{2}\xi_{y})$
$+4\pi G\rho_{0}^{2}\frac{k_{y}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (84)
$-\rho_{0}\omega^{2}\xi_{z}=$ $-[(3-2\beta)(k_{x}\xi_{x}+k_{y}\xi_{y})+3(k_{z}\xi_{z})]k_{z}p_{0}$
$+4\pi G\rho_{0}^{2}\frac{k_{z}}{k^{2}}(k_{x}\xi_{x}+k_{y}\xi_{y}+k_{z}\xi_{z})$ (85)
@@
$\xi_{\parallel}=\xi_{z},\quad\xi_{\perp}=(\xi_{x},\xi_{y}),\quad k_{\parallel}=k_{z},$ (86)
$\quad k_{\perp}=(k_{x},k_{y}),\quad k_{\perp}^{2}=k_{x}^{2}+k_{y}^{2}.$ (87)
We now multiply Eq. (B) by $k_{x}$ and Eq. (B) by $k_{y}$ and add the resulting equations together to yield
$\left[\rho\omega^{2}-(1+\beta)k_{\perp}^{2}p-\beta B^{2}k^{2}+4\pi G\rho^{2}\frac{k_{\perp}^{2}}{k^{2}}\right]k_{\perp}\cdot\xi_{\perp}$
$\qquad\qquad\qquad\qquad=\left(k_{\perp}^{2}p-4\pi G\rho^{2}\frac{k_{\perp}^{2}}{k^{2}}\right)k_{\parallel}\xi_{\parallel}$ (88)
Equation (B) yields
$\left(\rho\omega^{2}-3k_{\parallel}^{2}p+4\pi G\rho^{2}\frac{k_{\parallel}^{2}}{k^{2}}\right)k_{\parallel}\xi_{\parallel}$ (89)
$\qquad=\left((3-2\chi)k_{\parallel}^{2}p-4\pi G\rho^{2}\frac{k_{\parallel}^{2}}{k^{2}}\right)k_{\perp}\cdot\xi_{\perp}$ (90)